OpenStudy (anonymous):

What is the slope of the line 4x + 6y = 12?

OpenStudy (caozeyuan):

6y=12-4x, y=2-2/3x. So the slope is -2/3 medal please

OpenStudy (anonymous):

To get the slope, you have to rearrange the equation into y = mx + c form. Therefore. -6y = -4x + 12 You need a single y, so divide through by -6. y = -4x/-6 +12/-6 Your slope is the number in front of the x variable. I.e -4/-6 Simplify it to, 2/3